# Demonstration Organization via Logit Separability: Advancing In-Context Learning through Multiple Class-Related Words

### v4

@@ Abstract
Effective organization of in-context learning (ICL) demonstrations is key to improving the quality of large language model (LLM) responses. To create better demonstrations that guide LLM understanding, we introduce logit separability, a criterion to assess the clarity of both samples and class-related words at the logit level. This facilitates the optimization of sample and label selection, enhancing the precision of information provided in ICL demonstrations. Additionally, we find that incorporating multiple class-related words for each sample, rather than relying on a single class name, improves performance by offering a broader range of label information. Building on these insights, we propose LICL, a logit separability-based method that jointly organizes samples and integrates multiple class-related words into each sample-label pair. Evaluations across seven classification datasets show that this approach significantly improves ICL performance by providing clearer instructions and richer label information [^1] Our code is available at https://github.com/MidiyaZhu/MICL.
Demonstration Organization via Logit Separability: Advancing In-Context Learning through Multiple Class-Related Words
@@ 1 Introduction
Building on this concept, we observe that not all samples or class-related words exhibit the same degree of logit separability. As shown in Fig.1(a), some samples produce strong separation in the logits of class-related words, assigning high logits to words matching the true label and low logits to others, while others show negligible differences between them. Additionally, certain class-related words consistently display higher logit in samples of their corresponding label compared to samples of other labels. As shown in Fig.1(b), "bad" shows a clearer differentiation between positive and negative samples than the class name "negative".
Given the importance of providing clear label-instructed information in demonstrations, leveraging logit separability for selecting samples and labels might enhance ICL performance. Samples with strong logit separability provide more precise signals by aligning class-related words with the model’s internal representations (Fig.1(a)-accuracy). Furthermore, class-related words with distinct logit separability can outperform predefined class names in guiding the model toward accurate predictions. Therefore, selecting both samples and class-related words based on logit separability ensures clearer, more effective demonstrations, ultimately improving the performance of ICL.
@@
Given samples and labels’ critical roles in demonstration effectiveness for ICL, in this work, we propose a logit separability-based demonstration organization method. We organize demonstration samples based on their zero-shot logit separability across a pool of class-related words. In demonstration label selection, we propose a novel method that utilizes multiple class-related words instead of a single one. The selection, ordering, and quantity of class-related words are determined by the semantic correlations observed in the organized samples’ logit distribution and their validation performance. Moreover, while most research investigates either the sample selection (Zhang et al., 2022; Hongjin et al., 2022; Levy et al., 2023; Lu et al., 2022) or label selection (Yoo et al., 2022; Milios et al., 2023; Gao et al., 2023; Peskine et al., 2023) individually, we comprehensively address both aspects to enhance ICL quality.
@@ 3 Method
In this section, we introduce our proposed method LICL, which comprises logit-separability-based demonstrated sample selection and ordering, and corresponding multiple class-related words insertion, as shown in Fig.2.
@@ 3.1 Problem Statement
Give a large language model $M$, label space (class names) $L$, a class-related word pool $P$ with class-related words (includes class names), test sample $x_{test}$ and demonstrations $\left\{x_{i},y_{i}\right\}_{i=1}^{L}$ [^3] Our work focus on 1-shot ICL with label-balanced demonstrations, so labels in demonstration are from the $1$-th to the $L$-th label.. The zero-shot classification of $x_{test}$ can be based on the logit of class names as $\arg\max_{y\in L}p_{M}(y\mid x_{\text{test}})$ or based on the logit of all class-related words in the pool as $\arg\max_{y\in P}p_{M}(y\mid x_{\text{test}})$. Similarly, the 1-shot prediction of $x_{test}$ is $\arg\max_{y\in L}p_{M}(y\mid x_{1}\oplus y_{1},\cdots,x_{L}\oplus y_{L}\oplus x_{\text{test}})$ or $\arg\max_{y\in P}p_{M}(y\mid x_{1}\oplus y_{1},\cdots,x_{L}\oplus y_{L}\oplus x_{\text{test}})$. $\oplus$ is the concatenation operation, and function $p_{M}(\cdot)$ returns the logit of words in $M$’s vocabulary.
@@ 3.2 Logit Separability-Based Sample Selection and Ordering
Pool Refinement The initial pool $P$ is constructed from existing works Hu et al. (2022); Zhu et al. (2024) with details in Appx.D. However, not all words in $P$ exhibit clear separable features within both the dataset and the LLM, risking the introduction of incorrect class-related features. To refine $P$ and maintain the quality of the logit separability measurement, we propose a filtering method to generate a refined pool, $P_{r}$, by selecting class-related words based on their logit separability and correlation using point-biserial testing Kornbrot (2014) across dataset samples. Details of this refinement are in Appx.B, with analyses in Appx.H.1.
With the refined pool $P_{r}$ in place, for each zero-shot learning sample, the LLM generates logit values for all words in $P_{r}$. We first remove samples where the class-related word with the highest logit corresponds to an incorrect label. Next, we score samples for selection. Since predictions rely on the word with the highest logit, scoring based on all words introduces noise—particularly in multi-class tasks—because the pool contains far more words linked to incorrect labels than to the correct label. To address this, we focus exclusively on the true label pool, using two methods—Top-Logit Summation and Rank-Weighted Counting. These methods ensure that selected samples align with logit separability across relevant labels, yielding higher logits for class-related words matching the sample’s label and ranking them among the top in $P_{r}$. Let $N_{l}$ be the number of words in $P_{r}$ for class $l$, and $P_{r}^{l}$ be the pool of class-related words for class $l$.
@@ 3.3 Multiple Class-Related Words Insertion
After confirming that the selected samples respond well to these multiple class-related words in $P_{r}$, our method sequentially inserts the class-related words into a single sample-label pair to form a sample-multiple-label pair, which is more computationally and memory efficient than using multiple augmented sample-label pairs in ICL instruction.
@@ 4 Experiments
In this section, we examine the capacities of our method in ICL from the following perspectives: (1) 1-shot ICL classification comparison between baseline models and LICL (Sec.4.2); (2) 5-shot ICL (Sec.4.3); (3) Ablation studies on sample ordering (Sec.4.4.1) and multiple class-related words selection and ordering (Sec.4.4.2); (4) Scalability to larger LLMs (Sec.4.5), (5) General applicability of selected multiple class-related words (Sec.4.6); and (6) Further analyses (Sec.4.7).
@@ 4.1 Setups
Datasets We evaluate our method on seven datasets, including SST-2 (Socher et al., 2013), CR (Ding et al., 2008), IMDB (Maas et al., 2011), ISEAR (Scherer and Wallbott, 1994), AMAN (Aman and Szpakowicz, 2008), TREC-6 (Li and Roth, 2002) and AGNews Zhang et al. (2015). We adopt the templates in Wang et al. (2023). The initial class-related word pools are from Hu et al. (2022) and Zhu et al. (2024). Detailed information on the datasets, templates and label sets is given in Appx.D.
@@
For each model, multiple class-related words insertion is evaluated on validation set, which is the full training set excluding the samples selected for demonstrations. In Vanilla, TopK, and SelfICL, the full training set is split into a reduced training set and a validation set in an 8:2 ratio as demonstration samples vary with test samples. The demonstration samples are selected based on the reduced training set, while the multiple class-related words insertion is based on validation set. The ICL performance is evaluated on the same test set for all experiments.
@@ 4.2 Main Results
We present the results of each model and LICL (ours) under two settings: with and without multiple class-related words insertion in 1-shot ICL.
@@
+MLabels_CN 85.63$\uparrow$ 67.07$\uparrow$ 68.23$\uparrow$ 53.60$\uparrow$ 39.35$\uparrow$ 50.01$\uparrow$ 58.44$\uparrow$ 60.40
@@
+MLabels_CN 83.80$\uparrow$ 64.89 61.46$\uparrow$ 64.20$\uparrow$ 43.36$\uparrow$ 54.82$\uparrow$ 56.42$\uparrow$ 61.28
+MLabels_LW 83.80 64.89 62.16$\uparrow$ 65.40$\uparrow$ 43.11 55.08$\uparrow$ 56.42 61.55
@@
+MLabels_CN 84.84$\uparrow$ 63.83 69.80$\uparrow$ 58.20$\uparrow$ 36.84$\uparrow$ 48.04$\uparrow$ 51.14$\uparrow$ 58.96
+MLabels_LW 84.84 63.83 69.20 65.20$\uparrow$ 36.84 48.04 51.14 59.87
@@
Table 1: ICL Experimental Results: LICL is demonstrated on initial sample-label pairs while ‘+MLabels’ is on sample-multiple-label pairs. ‘+MLabels_CN’ refers to prediction based on the maximum logit over class names, and ‘+MLabels_LW’ refers to prediction based on the maximum logit over inserted class-related words. The best accuracy results (%) are marked in bold. Marker 1 indicates results given under Eq.1, while marker 2 indicates results given under Eq.2. Upward arrow ($\uparrow$) in ‘+MLabels_CN’ signifies an increase in performance compared to the original method, while in ‘+MLabels_LW’, it signifies an increase in performance compared to ‘+MLabels_CN’.
LICL Outperforms Over Baselines: As demonstrated in Table 1, the proposed LICL consistently outperforms the baseline methods across all datasets when evaluated under initial class-related word conditions. It achieves an average accuracy improvement of 3.11% in LLaMA2-7b and 11.58% in GPT2-xl. Particularly notable are the gains in multi-class classification tasks, where it records accuracy improvements of 6.02% (AMAN, LLaMA2-7b), 12.8% (TREC, GPT2-xl), 12.53% (AMAN, GPT2-xl), and 23.68% (AGNews, GPT2-xl).
Multiple Class-Related Words Enhanced: Utilizing multiple class-related words (‘+MLabels’) significantly improves performance across all models. This strategy results in an average baselines’ accuracy increase of over 2% in LLaMA2-7b, reaching up to 2.6% in SelfICL, and more than 5% in GPT2-xl, with a maximum improvement of 7.37% in vanilla GPT2-xl. Compared to configurations that employ a single class-related word per sample-label pair, LICL achieves notable accuracy gains, with increases of 2.33% in LLaMA2-7b (peaking at 9.6% in AMAN) and 3.84% in GPT2-xl (peaking at 11.09% in ISEAR). These improvements demonstrate significant enhancements over the initial single-label setup, underscoring the efficacy of the multiple class-related word approach.
Validation-Testing Results Alignment: Fig.3 displays the validation and test performance across various counts of inserted class-related words (N) for each dataset in LICL. Generally, the trends in test performance are consistent with those in validation performance. The peak test accuracy was achieved by incorporating multiple class-related words before reaching the stopping criterion, at which point the validation accuracy begins to decrease with further insertion of class-related words. This consistency supports the reliability of our stopping criterion.
@@
Initial Class-Related Word Updates Improved To improve the quality of the initial sample-label pairs, we jointly update certain labels in two datasets based on the samples selected in LICL and DataICL before ICL experiments [^4] Updating initial class-related words ensures they have a high logit value relative to other words with the same label polarity to provide effective class-related guidance in demonstrations for all testing samples. However, Vanilla, TopK, and SelfICL use varying demonstration samples across test instances, leading to inconsistent evaluations and making updates impractical. Therefore, we use class names as initial words in the multiple insertion experiment.. In AMAN, ‘other’ is replaced with ‘neutral’. In TREC, ‘entity’ is replaced with ‘animal’, ‘description’ with ‘definition’, ‘human’ with ‘persons’, ‘location’ with ‘state’, and ‘number’ with ‘numeric’. The enhanced accuracy presented in Table 2 highlights the effectiveness of class-related word updates over the direct use of predefined class names.
@@ 4.3 Effectiveness of LICL in 5-shot ICL
Table 3: Multiple class-related words insertion accuracy (%) in 5-shot ICL: Scoring method and test sets as in Table 1.
@@ 4.4.1 Effectiveness of LICL Ordering
In LICL, we use LLM’s feedback over class-related words as a decision feature to order sample-label pairs. To evaluate our ordering method’s effectiveness, we compare 1-shot ICL classification performance using LICL’s initial sample-label pairs against 30 (LLaMA2-7b) and 50 (GPT2-xl) random permutations (excluding LICL’s order) in multi-class tasks [^5] We randomize the order of samples selected by LICL to create various permutations and then randomly select 30 or 50 different permutations from the generated pool. Each permutation is assessed individually in a 1-shot ICL classification., and a flipped order in binary tasks.
@@
As shown in Table 4, LICL often outperforms or matches the best results among compared permutations. It achieves an average accuracy improvement of 0.39% in LLaMA2-7b and 5.86% in GPT2-xl. Despite a notable performance drop for IMDB (GPT2-xl), the binary task nature can mitigate this by enumerating all orders. LICL excels in multi-class tasks, matching or exceeding top permutation results. This highlights the effectiveness of LICL’s demonstration order, particularly in multi-class tasks with thousands of possible orderings ($7!$), which would take months to process enumerating in LLMs, whereas LICL achieves comparable performance within hours or minutes.

### v5

# Logit Separability-Driven Samples and Multiple Class-Related Words Selection for Advancing In-Context Learning

@@ Abstract
Effective organization of in-context learning (ICL) demonstrations is key to improving the quality of large language model (LLM) responses. To create better sample-label pairs that instruct LLM understanding, we introduce logit separability, a criterion to assess the clarity of both samples and class-related words at the logit level. This facilitates the optimization of sample and label selection, enhancing the precision of information provided in ICL demonstrations. Additionally, we find that incorporating multiple class-related words for each sample, rather than relying on a single class name, improves performance by offering a broader range of label information. Building on these insights, we propose LICL, a logit separability-based method that jointly organizes samples and integrates multiple class-related words into each sample-label pair. Evaluations across seven classification datasets show that this approach significantly improves ICL performance by providing clearer instructions and richer label information [^1] Our code is available at https://github.com/MidiyaZhu/MICL.
Logit Separability-Driven Samples and Multiple Class-Related Words Selection for Advancing In-Context Learning
@@ 1 Introduction
Building on this concept, we observe that not all samples or class-related words exhibit the same degree of logit separability. As shown in Fig.1(a), some samples produce strong separation in the logits of class-related words, assigning high logits to words matching the true label and low logits to others, while others show marginal differences between them. Additionally, certain class-related words consistently display higher logit in samples of their corresponding label compared to samples of other labels. As shown in Fig.1(b), "bad" shows a clearer differentiation between positive and negative samples than the class name "negative".
Given the importance of providing clear label-instructed information in demonstrations, leveraging logit separability for selecting samples and labels might enhance ICL performance. As shown in Fig.1(a)-accuracy, samples with strong logit separability provide more precise signals by aligning class-related words with the model’s internal representations. Besides, class-related words with distinct logit separability can beat predefined class names in guiding the model toward accurate predictions (Appx.A.2.1). Therefore, selecting both samples and class-related words via logit separability ensures clearer, more effective demonstrations, ultimately improving the performance of ICL.
@@
Given samples and labels’ critical roles in demonstration effectiveness for ICL, in this work, we propose a logit separability-based demonstration organization method. We organize demonstration samples based on their zero-shot logit separability across a pool of class-related words. In demonstration label selection, we propose a novel method that utilizes multiple class-related words instead of a single one. The selection, ordering, and quantity of class-related words are determined through sequential forward search, guided by logit feedback from the selected samples and their validation performance. Moreover, while most research investigates either the sample selection (Zhang et al., 2022; Hongjin et al., 2022; Levy et al., 2023; Lu et al., 2022) or label selection (Yoo et al., 2022; Milios et al., 2023; Gao et al., 2023; Peskine et al., 2023) individually, we comprehensively address both aspects to enhance ICL quality.
@@ 3 Method
In this section, we introduce our proposed method LICL, which comprises logit separability-based demonstrated sample selection and ordering, and corresponding multiple class-related words insertion, as shown in Fig.2.
@@ 3.1 Problem Statement
Give a large language model $M$, label space (class names) $L$, a class-related word pool $P$ with class-related words (includes class names), test sample $x_{test}$ and demonstrations $\left\{x_{i},y_{i}\right\}_{i=1}^{L}$ [^3] Our work focus on 1-shot ICL with label-balanced demonstrations, so labels in demonstration are from the $1$-th to the $L$-th label. Discussion of label balance can be found in Appx.H.4. The zero-shot classification of $x_{test}$ can be based on the logit of class names as $\arg\max_{y\in L}p_{M}(y\mid x_{\text{test}})$ or based on the logit of all class-related words in the pool as $\arg\max_{y\in P}p_{M}(y\mid x_{\text{test}})$. Similarly, the 1-shot prediction of $x_{test}$ is $\arg\max_{y\in L}p_{M}(y\mid x_{1}\oplus y_{1},\cdots,x_{L}\oplus y_{L}\oplus x_{\text{test}})$ or $\arg\max_{y\in P}p_{M}(y\mid x_{1}\oplus y_{1},\cdots,x_{L}\oplus y_{L}\oplus x_{\text{test}})$. $\oplus$ is the concatenation operation, and function $p_{M}(\cdot)$ returns the logit of words in $M$’s vocabulary.
@@ 3.2 Logit Separability-Based Sample Selection and Ordering
Pool Refinement The initial pool $P$ is constructed from existing works Hu et al. (2022); Zhu et al. (2024) with details in Appx.D. However, not all words in $P$ exhibit correct separable features within both the dataset and the LLM, risking the introduction of incorrect class-related features. To refine $P$ and maintain the quality of the logit separability measurement, we propose a filtering method to generate a refined pool, $P_{r}$, by selecting class-related words based on their logit separability and correlation using point-biserial testing Kornbrot (2014) across dataset samples. Details of this refinement are in Appx.B, with analyses in Appx.H.1.
With the refined pool $P_{r}$ in place, for each zero-shot learning sample, the LLM generates logit values for all words in $P_{r}$. We first remove samples where the class-related word with the highest logit corresponds to an incorrect label. Next, we score samples for selection. Since predictions only rely on the word with the highest logit, scoring based on all words introduces noise—particularly in multi-class tasks—because the pool contains far more words linked to incorrect labels than to the correct label. To address this, we focus exclusively on the true label pool, using two methods—Top-Logit Summation and Rank-Weighted Counting. These methods ensure that selected samples align with logit separability across relevant labels, yielding higher logits for class-related words matching the sample’s label and ranking them among the top in $P_{r}$. Let $N_{l}$ be the number of words in $P_{r}$ for class $l$, and $P_{r}^{l}$ be the pool of class-related words for class $l$.
@@ 3.3 Multiple Class-Related Words Insertion
After confirming that the selected samples respond well to the multiple class-related words in $P_{r}$, our method sequentially inserts these words into a single sample-label pair to form a sample-multiple-label pair, which is more computationally and memory efficient than using multiple augmented sample-label pairs in ICL instruction.
@@ 4 Experiments
In this section, we examine the capacities of our method in ICL from the following perspectives: (1) 1-shot ICL classification comparison between baseline models and LICL (Sec.4.2); (2) 5-shot ICL (Sec.4.3); (3) Ablation studies on sample ordering (Sec.4.4.1) and multiple class-related words selection and ordering (Sec.4.4.2); (4) Scalability to larger LLMs (Sec.4.5); (5) General applicability of selected multiple class-related words (Sec.4.6); and (6) Further analyses (Sec.4.7).
@@ 4.1 Setups
Datasets Seven datasets are evaluated, including: SST-2 (Socher et al., 2013), CR (Ding et al., 2008), IMDB (Maas et al., 2011), ISEAR (Scherer and Wallbott, 1994), AMAN (Aman and Szpakowicz, 2008), TREC-6 (Li and Roth, 2002), and AGNews (Zhang et al., 2015). The templates are provided by Wang et al. (2023). The initial class-related word pools are drawn from Hu et al. (2022) and Zhu et al. (2024). Detailed information on the datasets, templates, and pool is given in Appx.D.
@@
For each model, the sequential class-related word forward insertion is evaluated on the validation set, which is the full training set excluding the samples selected for demonstrations. In Vanilla, TopK, and SelfICL, the full training set is split into a reduced training set and a validation set in an 8:2 ratio as demonstration samples vary with test samples. The demonstration samples are selected based on the reduced training set, while the multiple class-related word insertion is based on the validation set. The ICL performance is evaluated on the same test set for all experiments.
@@ 4.2 Main Results
We present the results of each model and LICL (ours) under two settings: with and without multiple class-related word insertion in 1-shot ICL. The outcomes, summarized in Table 1, lead to the following conclusions:
@@
+MLabels_CN 85.63$\uparrow$ 67.07$\uparrow$ 68.23$\uparrow$ 53.60$\uparrow$ 39.35$\uparrow$ 50.01$\uparrow$ 58.44$\uparrow$ 60.33
@@
+MLabels_CN 83.80$\uparrow$ 65.43$\uparrow$ 61.46$\uparrow$ 64.20$\uparrow$ 43.36$\uparrow$ 54.82$\uparrow$ 56.42$\uparrow$ 61.36
+MLabels_LW 83.80 65.43 62.16$\uparrow$ 65.40$\uparrow$ 43.11 55.08$\uparrow$ 56.42 61.63
@@
+MLabels_CN 84.84$\uparrow$ 65.74$\uparrow$ 69.80$\uparrow$ 58.20$\uparrow$ 36.84$\uparrow$ 48.04$\uparrow$ 51.14$\uparrow$ 59.23
+MLabels_LW 84.84 65.74 69.20 65.20$\uparrow$ 36.84 48.04 51.14 60.14
@@
Table 1: 1-Shot ICL Experimental Results: Model Name (e.g., TopK, LICL) is demonstrated on initial sample-label pairs while ‘+MLabels’ is on sample-multiple-label pairs. ‘+MLabels_CN’ is predicting based on the maximum logit over class names, and ‘+MLabels_LW’ is based on the maximum logit over inserted class-related words. The best accuracy results (%) are marked in bold. Marker 1 indicates results given under Eq.1, while marker 2 indicates results given under Eq.2. $\uparrow$" in ‘+MLabels_CN’ signifies an increase in performance compared to the original method, while in ‘+MLabels_LW’, it signifies an increase in performance compared to ‘+MLabels_CN’.
LICL Outperforms Baselines: The proposed LICL consistently outperforms baseline methods across all datasets under the initial single-label pair conditions. It achieves an average accuracy improvement of 3.11% in LLaMA2-7b and 11.58% in GPT2-xl. The gains are especially significant in multi-class classification tasks, with notable improvements of 6.02% (AMAN, LLaMA2-7b), 12.80% (TREC, GPT2-xl), 12.53% (AMAN, GPT2-xl), and 23.68% (AGNews, GPT2-xl).
Multiple Class-Related Words Enhanced: Inserting multiple class-related words (‘+MLabels’) significantly improves performance across all models. This strategy yields an average baselines’ accuracy increase of 2.39% in LLaMA2-7b, reaching up to 2.6% in SelfICL, and an average increase of 5.94% in GPT2-xl, with a maximum improvement of 7.37% in vanilla GPT2-xl. Compared to configurations that employ a single class-related word per sample-label pair, LICL achieves notable accuracy gains, with increases of 2.33% in LLaMA2-7b (peaking at 9.6% in AMAN) and 3.84% in GPT2-xl (peaking at 11.09% in ISEAR). These improvements demonstrate significant enhancements over the initial single-label setup, underscoring the efficacy of the multiple class-related word approach.
Validation-Testing Results Alignment: Fig.3 displays the validation and test performance across various counts of inserted class-related words (N) for each dataset in LICL. Generally, the trends in test performance are consistent with those in validation performance. The peak test accuracy was achieved by incorporating multiple class-related words before reaching the stopping criterion, at which point the validation accuracy begins to decrease with further insertion of class-related words. This consistency supports the reliability of our stopping criterion in sequential forward insertion.
@@
Initial Class-Related Word Updates Improved To boost the quality of the initial sample-label pairs, we update certain labels in two datasets based on the samples selected in LICL and DataICL before ICL experiments [^4] Updating initial class-related words ensures they have a high logit value relative to other words with the same label polarity to provide effective class-related guidance in demonstrations for all testing samples. However, Vanilla, TopK, and SelfICL use varying demonstration samples across test instances, leading to inconsistent evaluations and making updates impractical. Therefore, we use class names as initial words in the multiple insertion experiment.. In AMAN, ‘other’ is replaced with ‘neutral’. In TREC, ‘entity’ is replaced with ‘animal’, ‘description’ with ‘definition’, ‘human’ with ‘persons’, ‘location’ with ‘state’, and ‘number’ with ‘numeric’. The enhanced accuracy presented in Table 2 highlights the effectiveness of class-related word updates over the direct use of predefined class names.
@@ 4.3 Effectiveness of LICL in 5-shot ICL
Table 3: Accuracy (%) with multiple class-related word insertion in 5-shot ICL: The scoring method and test sets are consistent with those in Table 1.
@@ 4.4.1 Effectiveness of LICL Ordering
In LICL, we use the logit separability scores of samples as a decision feature to order sample-label pairs. To assess the effectiveness of this ordering strategy, we compare 1-shot ICL classification performance using LICL’s initial sample-label pairs against 30 (LLaMA2-7b) and 50 (GPT2-xl) random permutations (excluding LICL’s order) in multi-class tasks [^5] We randomize the order of samples selected by LICL to create various permutations and then randomly select 30 or 50 different permutations from the generated pool. Each permutation is assessed individually in a 1-shot ICL classification., and a flipped order in binary tasks.
@@
As shown in Table 4, LICL often outperforms or matches the best results among compared permutations. It achieves an average accuracy improvement of 0.39% in LLaMA2-7b and 5.65% in GPT2-xl. Although GPT2-xl shows a performance drop on the IMDB dataset, the binary nature of this task can mitigate this by allowing all possible orders to be tested exhaustively. LICL excels in multi-class tasks, matching or exceeding top permutation results. This underscores LICL’s ordering strategy’s efficiency, especially in multi-class tasks with thousands of possible orderings ($7!$). While enumerating all these orderings could take months, LICL achieves comparable performance within hours or minutes, demonstrating its practical advantage.